Orthogonally diagonalize    as  

a.
b.
c.
d.
e.
f.

Orthogonally diagonalize A as PDPT A = [$ ] a. 1 2 P = 1 2 1 2 D [ 6 0 0-4 1 2 b . Р = 1 1 2 12 1 1 2 2 D = 16] -60 04 - C.
Answer #1

24м, а 44 и, то As Now 1A-A7 ( 0 (EA)2 - 25 го - А2 - - - -h=0 » (А - 6) (A+A) го А= 6-4 ¥0л. А= 6, AX 2GX Я, t Sl, - Gut г)Now pa 12 -1782 1/2 1/2 6 + Do ting (6-7) -(6-4) *HOW APA(47)C::) - 416) + PDet (1967) (673) Then Ap-PD => Ais onthogonally d